# Scalable Fine-tuning from Multiple Data Sources: A First-Order Approximation Approach

### v1

@@ Abstract
We study the problem of fine-tuning a language model (LM) for a target task by optimally using the information from $n$ auxiliary tasks. This problem has broad applications in NLP, such as targeted instruction tuning and data selection in chain-of-thought fine-tuning. The key challenge of this problem is that not all auxiliary tasks are useful to improve the performance of the target task. Thus, choosing the right subset of auxiliary tasks is crucial. Conventional subset selection methods, such as forward & backward selection, are unsuitable for LM fine-tuning because they require repeated training on subsets of auxiliary tasks. This paper introduces a new algorithm to estimate model fine-tuning performances without repeated training. Our algorithm first performs multitask training using the data of all the tasks to obtain a meta initialization. Then, we approximate the model fine-tuning loss of a subset using functional values and gradients from the meta initialization. Empirically, we find that this gradient-based approximation holds with remarkable accuracy for twelve transformer-based LMs. Thus, we can now estimate fine-tuning performances on CPUs within a few seconds. We conduct extensive experiments to validate our approach, delivering a speedup of 30×30$\times$30 × over conventional subset selection while incurring only $1\%$ error of the true fine-tuning performances. In downstream evaluations of instruction tuning and chain-of-thought fine-tuning, our approach improves over prior methods that utilize gradient or representation similarity for subset selection by up to $3.8\%$.
@@ 1 Introduction
Fine-tuning a language model (LM) has emerged as an effective approach for knowledge transfer on text data. As the scale of LMs keeps growing, efficient and scalable fine-tuning methods are in great demand. For instance, parameter-efficient fine-tuning with adapters or low-rank parameterization can significantly reduce the memory usage of fine-tuning Houlsby et al. (2019); Pfeiffer et al. (2020); Hu et al. (2021). In many applications, besides solving a target task of interest, one can access several related data sources, which could be used for data augmentation. Examples include multilingual systems such as neural machine translation Neubig and Hu (2018), parsing Üstün et al. (2020), data selection Xie et al. (2023), and targeted instruction tuning Xia et al. (2024). A fundamental issue in these scenarios is selecting the beneficial data sources amongst all the available data sources, which can be formulated as a subset selection problem. In this paper, we develop efficient algorithms for subset selection in LM fine-tuning given $n$ data sources, which can scale up to handle large numbers of $n$.
@@
Figure 1: An overview of our approach: (1) Perform multitask training on a base LM using all the samples, leading to a meta-initialization $\theta^{\star}\in{}^{p}$. We store (randomly) projected gradients of $\theta^{\star}$ for every sample. (2) Estimate model fine-tuning performances on a list of task subsets using projected gradients as features in logistic regression. (3) Using the estimated results (denoted as $\hat{f}(S_{1}),\dots,\hat{f}(S_{m})$), compute a score $T_{i}$ for each auxiliary task for $i=1,2,\dots,n$, which indicates its relevance to the target task. Select a subset (depicted in blue color) using a threshold score and combine their data for training jointly with the target task.
@@
Classical subset selection methods, such as forward & backward stepwise selection, have been very effective for regression analysis practice Hastie et al. (2009). However, these classical methods are unsuitable for LM fine-tuning because they require repeated training of the base LM on many subsets of tasks, which is not feasible for large $n$. More recently, data selection methods have been proposed based on selecting influential samples by comparing gradients Xia et al. (2024) or feature representation during training Ivison et al. (2023). These methods depend on the training procedure, leading to noisy outcomes for approximating true fine-tuning performances. In summary, existing subset selection methods require expensive computation or rely on noisy measurements of task-relatedness (Ruder, 2017; Vu et al., 2020, 2022; Zhang et al., 2023).
@@
This paper presents a new approach to scale up subset selection by quickly estimating model fine-tuning losses without running the fine-tuning procedure. The key idea is first to perform multitask training on the samples of all the tasks, leading to a meta initialization $\theta^{\star}$ Finn et al. (2017). From $\theta^{\star}$, we compute all the samples’ functional values and gradients, which can be used during inference. Second, given a subset of auxiliary tasks, we apply Taylor’s expansion on the loss and approximate the loss value with the function values and the gradients of the samples in the subset. After applying this approximation, we will estimate the fine-tuning losses with logistic regression. Notice that this computation can be done on CPUs, and we use random projections to reduce the dimension of the regression down to a few hundred. Using this procedure, we can estimate the fine-tuning performance for each subset in just a few seconds, which is significantly faster than full-model fine-tuning. Importantly, the accuracy of this approach hinges on the quality of the first-order approximation. We find that across twelve transformer-based LMs, including Llama-3-8B, the approximation error of the gradient-based expansion is at the order of $10^{-5}$ to $10^{-3}$, in regions close to $\theta^{\star}$. Based on this estimation, we can apply a classical subset selection method, such as forward selection, to every subset encountered during the procedure. In summary, our approach has three stages. See Figure 1 for an illustration.
@@
We conduct extensive experiments to validate our approach using four benchmarks. We show that GradEx can accurately approximate fine-tuned model losses within 1% errors across various LMs. Meanwhile, GradEx can accelerate forward selection with 30$\times$ fewer FLOPs and 25$\times$ less GPU hours by circumventing full model fine-tuning. GradEx also speeds up random ensemble with 44$\times$ fewer FLOPs and 46$\times$ less GPU hours. We consider targeted instruction tuning and chain-of-thought fine-tuning for downstream evaluations of GradEx. Our approach matches the accuracy of classical subset selection methods with full fine-tuning while costing only 0.5% of the computation. On the ToxiGen and TruthfulQA benchmarks, GradEx outperforms existing selection methods that rely on gradient or representation measurements Xie et al. (2023); Xia et al. (2024) by 3.8% on average. On StrategyQA and CommonsenseQA datasets, GradEx improves chain-of-thought reasoning accuracy by 2.4% over prior methods.
@@
In conclusion, this paper introduces a new algorithm for subset selection in LM fine-tuning given multiple data sources. The algorithm can quickly estimate model fine-tuning performances without performing fine-tuning on each subset. We identify a linearization property of LMs (after meta-training), which is empirically verified on twelve transformer-based LMs. This empirical finding may be relevant for future developments of fast inference methods of LLMs. Our estimation algorithm unlocks classical subset selection methods such as forward & backward selection to LMs. Our extensive experiments show that our approach can deliver significant speedups without losing downstream accuracy. In particular, our approach outperforms recent subset selection methods for instruction tuning and chain-of-thought fine-tuning. The code repository and detailed instructions for reproducing our findings are available at https://github.com/VirtuosoResearch/Scalable-finetuning.
@@ 2.1 Prior approaches for measuring $f(S)$
Experimental setup: We will present a case study using the Alpaca dataset Taori et al. (2023), which involves $52,000$ instruction-following data generated by GPT-3, each containing an instruction as a task description and the corresponding task input and output. The dataset can be pre-processed into $38$ tasks based on the instruction types Wang et al. (2023b). Each task is identified by the verb in the instruction, such as “edit” and “describe.” We randomly sample $10\%$ as the validation set. We use TinyLlama-1.1B as the base LM and LoRA as the base protocol Hu et al. (2021).
@@
$\diamondsuit$ Third, we also find that methods that utilize model gradients Fifty et al. (2021) or representation similarity Wu et al. (2020); Vu et al. (2020) to select tasks do not correlate well with actual values of $f$. For every pair of tasks $i$ and $j$, we evaluate the correlation between the cosine similarity of the averaged gradient of tasks $i,j$, and $f(\{i,j\})$. We find it is less than $0.2$.
@@ 2.2 Problem statement
Having described the nuances of $f$ and the potential complexity in optimizing $f$ directly (see also Zhang et al. (2023) for further discussion on this topic), we now pose the following research question: Given a list of subsets $S_{1},\dots,S_{m}\subseteq\{1,\ldots,n\}$, can we efficiently estimate $f(S_{1}),\dots,f(S_{m})$, without fine-tuning the model on each possible subset? Next, we give two examples to illustrate the choices of subsets.
@@
If we can quickly estimate the values of $f(S_{1}),f(S_{2}),\dots,f(S_{m})$, then we can still apply the above subset selection methods. In the next section, we present our estimation approach.
@@ 3 Our Approach
Table 1: We empirically find that the first-order approximation holds with very high accuracy within 10% relative distance to the meta initialization $\theta^{\star}$. As shown in the table below, the RRSS is at the order of $10^{-5}$-$10^{-3}$ when $X$ is between $2\%$-$10\%$ distance of $\theta^{\star}$. We attribute this behavior to the highly overparameterized nature of LLMs. We report the average over $50$ random task subsets to ensure statistical significance. The reference for each LM can be found in Appendix D.
Dist. Pythia-70M BERT-Base RoBERTa-Base GPT-2 FLAN-T5-Base BloomZ-560M
2% $9_{\pm 1.4}\times 10^{-4}$ $3_{\pm 0.2}\times 10^{-4}$ $4_{\pm 0.4}\times 10^{-4}$ $2_{\pm 0.2}\times 10^{-4}$ $1_{\pm 0.2}\times 10^{-4}$ $2_{\pm 0.4}\times 10^{-4}$
4% $1_{\pm 0.2}\times 10^{-4}$ $5_{\pm 0.9}\times 10^{-4}$ $5_{\pm 0.5}\times 10^{-4}$ $6_{\pm 0.7}\times 10^{-4}$ $8_{\pm 0.4}\times 10^{-4}$ $5_{\pm 1.5}\times 10^{-4}$
6% $3_{\pm 0.3}\times 10^{-4}$ $9_{\pm 1.0}\times 10^{-4}$ $6_{\pm 0.9}\times 10^{-4}$ $8_{\pm 0.6}\times 10^{-4}$ $2_{\pm 0.3}\times 10^{-4}$ $9_{\pm 0.5}\times 10^{-4}$
8% $4_{\pm 0.9}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-3}$ $9_{\pm 1.3}\times 10^{-4}$ $3_{\pm 0.3}\times 10^{-3}$ $3_{\pm 0.6}\times 10^{-4}$ $7_{\pm 0.6}\times 10^{-4}$
10% $7_{\pm 1.4}\times 10^{-3}$ $5_{\pm 1.4}\times 10^{-3}$ $5_{\pm 0.5}\times 10^{-3}$ $5_{\pm 0.4}\times 10^{-3}$ $5_{\pm 1.2}\times 10^{-3}$ $5_{\pm 2.2}\times 10^{-3}$
Dist. TinyLlama-1.1B GPT-Neo-1.3B OPT-1.3B Gemma-2-2B Mistral-7B Llama-3-8B
2% $6_{\pm 0.5}\times 10^{-5}$ $3_{\pm 0.4}\times 10^{-5}$ $7_{\pm 0.2}\times 10^{-5}$ $4_{\pm 0.3}\times 10^{-5}$ $9_{\pm 0.5}\times 10^{-5}$ $3_{\pm 0.3}\times 10^{-5}$
4% $1_{\pm 0.3}\times 10^{-4}$ $3_{\pm 0.3}\times 10^{-4}$ $7_{\pm 1.0}\times 10^{-5}$ $2_{\pm 0.1}\times 10^{-4}$ $2_{\pm 0.2}\times 10^{-4}$ $6_{\pm 0.9}\times 10^{-5}$
6% $3_{\pm 0.7}\times 10^{-4}$ $6_{\pm 0.6}\times 10^{-4}$ $8_{\pm 0.2}\times 10^{-5}$ $3_{\pm 0.4}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-4}$ $3_{\pm 0.6}\times 10^{-4}$
8% $4_{\pm 0.9}\times 10^{-4}$ $1_{\pm 0.2}\times 10^{-3}$ $1_{\pm 0.1}\times 10^{-4}$ $8_{\pm 0.7}\times 10^{-4}$ $4_{\pm 0.2}\times 10^{-4}$ $4_{\pm 0.4}\times 10^{-4}$
10% $5_{\pm 0.8}\times 10^{-3}$ $5_{\pm 0.8}\times 10^{-3}$ $5_{\pm 0.1}\times 10^{-3}$ $4_{\pm 0.5}\times 10^{-3}$ $4_{\pm 0.1}\times 10^{-3}$ $5_{\pm 0.4}\times 10^{-4}$
@@ 3.1 Multitask training on all tasks
Let the output of a network be $h_{X}(s,y)$, where $s$ is an input (e.g., sentence) and $y$ is the prediction label. For example, in binary classification, $h_{X}$ is the log loss. Here, $X\in{}^{p}$ denotes the trainable parameters of all layers, while $\theta^{\star}\in{}^{p}$ denotes trained parameters.
@@
Table 1 reports the results, averaged over $50$ randomly sampled subsets with fixed sizes (3 for GLUE and 19 for Alpaca). Remarkably, across a range of relative distance values (between $X$ and $\theta^{\star}$) from 2% to 10%, the RRSS is at the range of $10^{-5}$ to $10^{-3}$.

### v2

@@ Abstract
We study the problem of fine-tuning a language model (LM) for a target task by optimally using the information from $n$ auxiliary tasks. This problem has broad applications in NLP, such as targeted instruction tuning and data selection in chain-of-thought fine-tuning. The key challenge of this problem is that not all auxiliary tasks are useful to improve the performance of the target task. Thus, choosing the right subset of auxiliary tasks is crucial. Conventional subset selection methods, such as forward and backward stepwise selection, are unsuitable for LM fine-tuning because they require repeated training on subsets of auxiliary tasks. This paper introduces a new algorithm to estimate model fine-tuning performances without repeated training. Our algorithm first performs multitask training using the data of all the tasks to obtain a meta initialization. Then, we approximate the model fine-tuning loss of a subset using functional values and gradients from the meta initialization. Empirically, we find that this gradient-based approximation holds with remarkable accuracy for twelve transformer-based LMs. Thus, we can now estimate fine-tuning performances on CPUs within a few seconds. Finally, we fine-tune the pretrained base model for once on the selected subset of tasks. We conduct extensive experiments to validate this approach, delivering a speedup of 30×30$\times$30 × over conventional subset selection while incurring only $1\%$ error of the true fine-tuning performances. In downstream evaluations involving both instruction tuning and chain-of-thought fine-tuning, this loss-based selection approach improves over prior gradient or representation similarity-based methods for subset selection by up to $3.8\%$.
@@ 1 Introduction
Fine-tuning a language model (LM) has emerged as an effective approach for knowledge transfer on text data. As the scale of LMs keeps growing, efficient and scalable fine-tuning methods are in great demand. For instance, parameter-efficient fine-tuning with adapters or low-rank parameterization can significantly reduce the memory usage of fine-tuning on a large model Houlsby et al. (2019); Pfeiffer et al. (2020); Hu et al. (2021). In many applications, besides solving a target task of interest, one can also access several related data sources, which could be used for data augmentation. Examples include multilingual systems such as neural machine translation Neubig and Hu (2018), parsing Üstün et al. (2020), data selection Xie et al. (2023), and targeted instruction tuning Xia et al. (2024). A fundamental issue in these scenarios is selecting the beneficial data sources amongst all the available data sources, which can be formulated as a subset selection problem. In this paper, we develop efficient algorithms for subset selection in LM fine-tuning given $n$ data sources, which can scale up to handle large numbers of $n$.
@@
Figure 1: An overview of our approach: (1) Perform multitask training on a base LM using all the samples, leading to a meta-initialization $\theta^{\star}\in{}^{p}$. We store (randomly) projected gradients of $\theta^{\star}$ for every sample. (2) Estimate model fine-tuning performances on a list of task subsets using projected gradients as features in logistic regression. (3) Using the estimated results (denoted as $\hat{f}(S_{1}),\dots,\hat{f}(S_{m})$), compute a score $T_{i}$ for each auxiliary task for $i=1,2,\dots,n$, which indicates its relevance to the target task. Finally, select a subset (depicted in blue color) using a threshold score, and fine-tune a pretrained LM by combining the training data of the selected tasks together.
@@
Classical subset selection methods, such as forward and backward stepwise selection, have been very effective for feature selection in regression analysis Hastie et al. (2009). However, these classical methods are unsuitable for LM fine-tuning because they require repeated training of the base LM on many subsets of tasks, which is not feasible for large $n$. More recently, data selection methods have been proposed based on selecting influential samples by comparing gradients Xia et al. (2024) or feature representation during training Ivison et al. (2023). These methods depend on the training procedure, leading to noisy outcomes for approximating true fine-tuning performances. In summary, existing subset selection methods require expensive computation or rely on noisy measurements of task-relatedness (Ruder, 2017; Vu et al., 2020, 2022; Zhang et al., 2023).
@@
This paper presents a new approach to scale up subset selection by quickly estimating model fine-tuning losses without running the fine-tuning procedure. The key idea is first to perform multitask training on the samples of all the tasks, leading to a meta initialization $\theta^{\star}$. From $\theta^{\star}$, we compute all the samples’ functional values and gradients, which can be used during inference. Second, given a subset of auxiliary tasks, we apply Taylor’s expansion on the loss and approximate the loss value with the function values and the gradients of the samples in the subset. After applying this approximation, we will estimate the fine-tuning losses directly with logistic regression without conducting any actual fine-tuning. Notice that this computation can be done on CPUs, and we use random projections to reduce the dimension of the regression down to a few hundred. Using this procedure, we can estimate the fine-tuning performance for each subset in just a few seconds, which is significantly faster than full-model fine-tuning. Importantly, the accuracy of this approach hinges on the quality of the first-order approximation. We find that across twelve transformer-based LMs, including Llama-3-8B, the approximation error of the gradient-based expansion is at the order of $10^{-5}$ to $10^{-3}$, in regions close to $\theta^{\star}$. Based on this estimation, we can apply a classical subset selection method, such as forward selection, to estimate the loss values of every subset encountered during the procedure. Lastly, we fine-tune a pretrained LM by combining the data from the selected subset of tasks. In summary, our approach has three stages. See Figure 1 for an illustration of our overall approach.
@@
We conduct extensive experiments to validate our approach using four benchmarks. We show that GradEx can accurately approximate fine-tuned model losses within 1% errors across various LMs. Meanwhile, GradEx can accelerate forward selection with 30$\times$ fewer FLOPs and 25$\times$ less GPU hours by circumventing full model fine-tuning. GradEx also speeds up random ensemble with 44$\times$ fewer FLOPs and 46$\times$ less GPU hours. We consider targeted instruction tuning and chain-of-thought fine-tuning for downstream evaluations of GradEx. Our approach matches the accuracy of classical subset selection methods with full fine-tuning while costing only 0.5% of the computation. On the ToxiGen and TruthfulQA benchmarks, GradEx outperforms existing selection methods that rely on gradient or representation measurements Xie et al. (2023); Xia et al. (2024) by 3.8% on average. On StrategyQA and CommonsenseQA datasets, the loss-based selection of GradEx improves chain-of-thought reasoning accuracy by 2.4% over gradient-based and feature-based selection methods.
@@
In conclusion, this paper introduces a new algorithm for subset selection in LM fine-tuning given multiple data sources. The algorithm can quickly estimate model fine-tuning performances without performing fine-tuning on each subset. We identify a linearization property of LMs (after meta-training), which is empirically verified on twelve transformer-based LMs. This empirical finding may be relevant for future developments of fast inference methods of LLMs. Our estimation algorithm unlocks classical subset selection methods such as forward & backward selection to LMs. The extensive experiments show that our approach can deliver significant speedups without losing downstream accuracy. In particular, our loss-based approach can outperform existing gradient-based or feature-based selection methods for instruction tuning and chain-of-thought fine-tuning. The code repository and detailed instructions for reproducing our findings are available at https://github.com/VirtuosoResearch/Scalable-finetuning.
@@ 2.1 Prior approaches for measuring $f(S)$
Experimental setup: We will present a case study using the Alpaca dataset Taori et al. (2023), which involves $52,000$ instruction-following data generated by GPT-3, each containing an instruction as a task description and the corresponding task input and output. The dataset can be pre-processed into $38$ tasks based on the instruction types Wang et al. (2023b). Each task is identified by the verb in the instruction, such as “edit” and “describe.” We randomly sample $10\%$ as the validation set and use the rest of $90\%$ samples as the training set. We use TinyLlama-1.1B as the base LM and LoRA as the base protocol Hu et al. (2021).
@@
$\diamondsuit$ Third, we also find that methods that utilize model gradients Fifty et al. (2021) or representation similarity Wu et al. (2020); Vu et al. (2020) to select tasks do not correlate well with actual values of $f$. For every pair of tasks $i$ and $j$, we evaluate the correlation between the cosine similarity of the averaged gradient of tasks $i,j$, and $f(\{i,j\})$. We find that this is less than $0.2$.
@@ 2.2 Problem statement
Having described the nuances of $f$ and the potential complexity in optimizing $f$ directly (see also Zhang et al. (2023) for further discussion on this topic), we now pose the following research question: Given a list of subsets $S_{1},\dots,S_{m}\subseteq\{1,\ldots,n\}$, can we efficiently estimate $f(S_{1}),\dots,f(S_{m})$, without fine-tuning the model on each possible subset? Next, we give two examples to illustrate the choices of these $m$ subsets.
@@
If we can quickly estimate the values of $f(S_{1}),f(S_{2}),\dots,f(S_{m})$, then we can still apply the above subset selection methods. In the next section, we present our estimation approach to obtain these loss values.
@@ 3 Our Approach
Table 1: We empirically find that the first-order approximation holds with very high accuracy within 0.25% relative distance to the meta initialization $\theta^{\star}$. As shown in the table below, the RRSS is at the order of $10^{-5}$-$10^{-3}$ when $X$ is between $0.05\%$ to $0.25\%$ distance of $\theta^{\star}$, where $\theta^{\star}$ includes all model weights in the initialization. We attribute this behavior to the highly overparameterized nature of LLMs. As a remark, we use LoRA as the base fine-tuning behavior; thus, the fine-tuned distance is small because of using LoRA. We report the average over $50$ random task subsets to ensure statistical significance. The reference for each LM can be found in Appendix D.
Distance Pythia-70M BERT-Base RoBERTa-Base GPT-2 FLAN-T5-Base BloomZ-560M
0.05% $9_{\pm 1.4}\times 10^{-4}$ $3_{\pm 0.2}\times 10^{-4}$ $4_{\pm 0.4}\times 10^{-4}$ $2_{\pm 0.2}\times 10^{-4}$ $1_{\pm 0.2}\times 10^{-4}$ $2_{\pm 0.4}\times 10^{-4}$
0.10% $1_{\pm 0.2}\times 10^{-4}$ $5_{\pm 0.9}\times 10^{-4}$ $5_{\pm 0.5}\times 10^{-4}$ $6_{\pm 0.7}\times 10^{-4}$ $8_{\pm 0.4}\times 10^{-4}$ $5_{\pm 1.5}\times 10^{-4}$
0.15% $3_{\pm 0.3}\times 10^{-4}$ $9_{\pm 1.0}\times 10^{-4}$ $6_{\pm 0.9}\times 10^{-4}$ $8_{\pm 0.6}\times 10^{-4}$ $2_{\pm 0.3}\times 10^{-4}$ $9_{\pm 0.5}\times 10^{-4}$
0.20% $4_{\pm 0.9}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-3}$ $9_{\pm 1.3}\times 10^{-4}$ $3_{\pm 0.3}\times 10^{-3}$ $3_{\pm 0.6}\times 10^{-4}$ $7_{\pm 0.6}\times 10^{-4}$
0.25% $7_{\pm 1.4}\times 10^{-3}$ $5_{\pm 1.4}\times 10^{-3}$ $5_{\pm 0.5}\times 10^{-3}$ $5_{\pm 0.4}\times 10^{-3}$ $5_{\pm 1.2}\times 10^{-3}$ $5_{\pm 2.2}\times 10^{-3}$
Distance TinyLlama-1.1B GPT-Neo-1.3B OPT-1.3B Gemma-2-2B Mistral-7B Llama-3-8B
0.05% $6_{\pm 0.5}\times 10^{-5}$ $3_{\pm 0.4}\times 10^{-5}$ $7_{\pm 0.2}\times 10^{-5}$ $4_{\pm 0.3}\times 10^{-5}$ $9_{\pm 0.5}\times 10^{-5}$ $3_{\pm 0.3}\times 10^{-5}$
0.10% $1_{\pm 0.3}\times 10^{-4}$ $3_{\pm 0.3}\times 10^{-4}$ $7_{\pm 1.0}\times 10^{-5}$ $2_{\pm 0.1}\times 10^{-4}$ $2_{\pm 0.2}\times 10^{-4}$ $6_{\pm 0.9}\times 10^{-5}$
0.15% $3_{\pm 0.7}\times 10^{-4}$ $6_{\pm 0.6}\times 10^{-4}$ $8_{\pm 0.2}\times 10^{-5}$ $3_{\pm 0.4}\times 10^{-4}$ $3_{\pm 0.4}\times 10^{-4}$ $3_{\pm 0.6}\times 10^{-4}$
0.20% $4_{\pm 0.9}\times 10^{-4}$ $1_{\pm 0.2}\times 10^{-3}$ $1_{\pm 0.1}\times 10^{-4}$ $8_{\pm 0.7}\times 10^{-4}$ $4_{\pm 0.2}\times 10^{-4}$ $4_{\pm 0.4}\times 10^{-4}$
0.25% $5_{\pm 0.8}\times 10^{-3}$ $5_{\pm 0.8}\times 10^{-3}$ $5_{\pm 0.1}\times 10^{-3}$ $4_{\pm 0.5}\times 10^{-3}$ $4_{\pm 0.1}\times 10^{-3}$ $5_{\pm 0.4}\times 10^{-4}$
@@ 3.1 Multitask training on all tasks
Let the output of a network be $h_{X}(s,y)$, where $s$ is an input (e.g., sentence) and $y$ is the prediction label. For example, in binary classification, $h_{X}$ is the log loss. Here, $X\in{}^{p}$ denotes the trainable parameters of all layers, while $\theta^{\star}\in{}^{p}$ denotes trained parameters. We obtain a model initialization $\theta^{\star}$ by multitask training on all tasks.
@@
We first verify that the fine-tuned models on subsets of tasks stay close to the initialization. We fine-tune one LM on a random subset of tasks until the validation loss stops decreasing. We measure the relative distance from the fine-tuned model weights to the initialization, i.e., $\frac{\|X-\theta^{\star}\|_{{}_{F}}}{\|\theta^{\star}\|_{{}_{F}}}$ over 50 task subsets. We consider all model weights in the initialization as $\theta^{\star}$. For full model fine-tuning, the relative distance remains within 0.25% on average. For LoRA fine-tuning, the relative distance stays within 0.19%.
Then, we report the first-order approximation error in Table 1. The results are averaged over $50$ randomly sampled subsets with fixed sizes (3 for GLUE and 19 for Alpaca). Remarkably, across a range of relative distance values (between $X$ and $\theta^{\star}$) from 0.05% to 0.25%, the RRSS is at the range of $10^{-5}$ to $10^{-3}$.